OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −6(2u − 1)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

-6 . (2u) -6 . (-1) Just that :)

OpenStudy (anonymous):

do -6 x 2u and -6 * -1

OpenStudy (anonymous):

@katiebell14:Got it ?

OpenStudy (skullpatrol):

@E.ali please don't do all the work for the student.

OpenStudy (anonymous):

@skullpatrol : So what I should to say !!

OpenStudy (anonymous):

I don't get how to do it

OpenStudy (anonymous):

Wait ..:)

OpenStudy (anonymous):

answer it : 2(6+5)

OpenStudy (anonymous):

@katiebell14 Take -6 and multiply it by the 2 numbers in the ( )

OpenStudy (anonymous):

or just give him the answer.................

OpenStudy (skullpatrol):

$$\huge a(b+c)=ab+ac$$

OpenStudy (skullpatrol):

The "a" is "distributed" as multiplier.

OpenStudy (anonymous):

Give a=-6 2u=b -1=c

OpenStudy (skullpatrol):